In a survey of 600 adults, 450 reported that e-mails are easy to misinterpret.

A. Construct a 99% confidence interval estimate for the population proportion of adults who report that e-mails are easy to misinterpret.

a. [0.601 , 0.816]
b. [0.704 , 0.796]
c. [0.729 , 0.739]
d. [0.684 , 0.895]
e. [0.736 , 0.785]
Answer #1

Solution :

Given that,

n = 600

x = 450

\hat p = x / n = 450 / 600 = 0.750

1 - \hat p = 1 - 0.750 = 0.250

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 2.576 * (\sqrt((0.750 * 0.250) / 600)

= 0.046

A 99% confidence interval for population proportion p is ,

\hat p - E < P <\hat p + E

0.750 - 0.046 < p < 0.750 + 0.046

0.704 < p < 0.796

(0.704 , 0.796)

Option b) is correct .